# STAGNet: A Spatio-Temporal Graph and LSTM Framework for Accident Anticipation

### v1

@@ 4.1.2 Networks and Baselines
We employ the pretrained object detection model YOLO [18] to extract object-level features. Specifically, we utilize the top 19 detected objects, each represented by a feature vector of dimension $d_{1}=4096$. To encode object labels, we apply GloVe word embeddings [19] using the spaCy library, resulting in embedding vectors of size $d_{2}=300$. Since the video clips are short (5 seconds in length), we set the temporal edge window $u=1$, meaning each object node is connected only to its counterpart in the previous frame. We fix the number of temporal neighbors at $k=20$. The architecture and configuration used across all datasets are consistent and illustrated in Fig. 2.
For global visual feature extraction, we adopt the SlowFast network [9], which produces a $h$-dimensional feature vector for each frame, where $h=2304$. All experiments are implemented using the PyTorch framework.
On the DAD dataset, we compare our approach against several accident anticipation methods: DSA [8], adaLEA [12], UString [5], DSTA [15], EGSTL [3], and Graph(Graph) [6]. These algorithms were selected based on their high relevance to this research and widespread use as comparison methods in similar work. The results for these comparative methods are directly taken from the respective publications. We used the same training/testing splits of DAD used by others to obtain the results for our proposed STAGNet model.
On the DoTA and DADA datasets, we compared our model against UString [5], UString (Geometric), and Graph(Graph) [6]. We were unable to compare with the rest of the methods as their implementations were not available for us to run them on DoTA and DADA datasets. Note that we implemented UString (Geometric), as an attempt to mimic GSTL [3], by replacing VGG-based object-level features of UString [5] with geometric features of the detected bounding boxes. Furthermore, although Mahmood et al. [3] have used a similar number of videos from the DoTA dataset, the subset of accident videos selected is different from what we have selected as we specifically selected ego-involved accidents, and they did not. Therefore, results reported in [3] on DoTA are not comparable to what we have achieved.
@@ 4.1.3 Evaluation Metrics
Accident anticipation models are expected to be both accurate and early in their predictions. Consistent with prior works [5, 3, 6, 1], we evaluate our model using Average Precision (AP) and mean Time-to-Accident (mTTA). AP reflects the correctness of predictions by measuring how well the model distinguishes frames preceding an accident from those with no accident in the near future. For a given frame, if the predicted probability $p_{t}$ of a future accident exceeds a threshold $\alpha$, it is counted as a positive prediction. By varying $\alpha$, a precision-recall curve is generated, and AP is computed as the area under this curve. A higher AP indicates more accurate predictions.
To determine how early the model can anticipate an accident, we use the Time-to-Accident (TTA) metric, which measures the time difference between the actual onset of an accident and the earliest frame, where $p_{t}>\alpha$. By evaluating TTA across various thresholds $\alpha$, we compute the mean TTA (mTTA) as the average over all corresponding values.
To reduce bias introduced by data splitting, we perform 5-fold cross-validation, and report the mean of the AP and mTTA values over all folds.
@@ 4.2.1 Comparison with State-of-the-art Methods
Table 2 shows the results of the experiments on the DAD dataset. Our proposed algorithm achieves the highest AP and mTTA, closely followed by Graph(Graph) [6]. These two methods perform considerably better than the rest of the comparison methods. The transformer-based method, AAT-ODA [24] achieves very competitive AP but at the expense of a lower mTTA.
Table 2: Performance comparison on DAD Dataset
@@
DAD  [8] DSA [8] 48.10 1.34
adaLEA [12] 52.30 3.43
UString [5] 53.70 3.53
DSTA [15] 56.1 3.66
EGSTL [3] 56.30 3.23
Graph(Graph)  [6] 62.80 4.29
AAT-DA  [24] 64.0 2.87
STAGNet (Ours) 64.36 4.54
As shown in Table 3, for DoTA and DADA datasets with ego-involved accidents, our model clearly outperforms all comparison methods across both datasets. On the DoTA dataset, the UString (Geometrics) and UString models achieve average precisions of 63.16% and 75.40%, respectively, while the Graph(Graph) model reaches 83.50%. In comparison, our model achieves an average precision of 92.73% and an mTTA of 3.23 seconds, meaning the model predicts accidents earlier and more reliably. On the DADA dataset, the performance gap is even more pronounced: our model achieves 95.39% AP, outperforming UString (Geometrics) at 56.75%, UString at 63.95%, and Graph(Graph) at 84.04%. The corresponding mTTA of 3.08s also confirms our methods’ ability to anticipate accidents earlier. These consistent improvements over multiple datasets highlight the effectiveness and robustness of our proposed framework in capturing rich spatio-temporal patterns for early and accurate accident anticipation. In particular, our improved frame-level feature pipeline has a marked effect on predicting ego-involved accidents.
@@
DoTA  [22] UString (Geometrics)  [3] 63.16 2.14
UString  [5] 75.40 2.26
Graph(Graph) [6] 83.50 3.14
@@
DADA  [23] UString (Geometrics)  [3] 56.75 1.25
UString  [5] 63.95 1.03
Graph(Graph)  [6] 84.04 2.92
@@ 4.2.2 Cross-dataset Evaluation
To assess the generalization capability of our model, we conducted cross-dataset experiments by training on one dataset and evaluating on another. As shown in Table 4, our method consistently outperforms the comparison methods across both datasets. When trained on the DoTA dataset and tested on DADA, our method achieves an AP of 67.47%, outperforming UString (Geometrics) at 46.39%, UString at 48.26%, and Graph(Graph) at 61.34%. Additionally, it achieves a higher mTTA of 3.28s, indicating better anticipation of accidents. Similarly, when trained on DADA and evaluated on DoTA, our method reaches an AP of 80.15%, surpassing UString (Geometrics) at 54.85%, UString at 55.63%, and Graph at 68.55%. It also records a higher mTTA of 3.13s. These results highlight the better generalization ability of our model across datasets.
However, the cross-dataset performance of all tested algorithms is considerably worse than cross-validation results on the same datasets. This can be attributed to the differences in video quality between the two datasets, particularly the higher bit rate  [3] in DoTA compared to DADA. The higher bit rate in DoTA results in clearer object boundaries and smoother motion, which can enable the model to learn more detailed spatial-temporal features. However, these learned features may not generalize well to the lower-quality videos in DADA, which contain more compression  [^2] https://github.com/JWFangit/LOTVS-DADA artifacts and reduced detail, thereby degrading cross-dataset performance.
@@
DoTA  [22] DADA (bAP = 40.8) UString (Geometrics)  [3] 46.39 1.71
UString  [5] 48.26 1.95
Graph(Graph) [6] 61.34 3.12
@@
DADA  [23] DoTA (bAP = 50.0) UString (Geometrics)  [3] 54.85 2.45
UString  [5] 55.63 2.17
Graph(Graph) [6] 68.55 2.86
@@
We analyze the impact of various global feature extraction methods on our accident anticipation framework. Since spatial features extracted from pretrained image-based networks such as VGG-16 focus solely on static appearance, they lack the ability to capture temporal dynamics essential for anticipating accidents. To overcome this limitation, spatio-temporal features extracted using I3D and SlowFast networks were explored. As shown in Table 5, the I3D model improves performance over VGG-16 by incorporating motion information across frames. Furthermore, our proposed use of SlowFast yields the best performance, outperforming both VGG16 and I3D. This improvement is attributed to SlowFast’s dual-pathway design, which effectively captures both fast and slow motion cues, providing a richer and more comprehensive spatio-temporal representation. These capabilities make SlowFast highly effective for predicting future accidents.
Table 5: Comparison of the effects of different Global Feature Extractors on our STAGNet model
@@
DoTA  [22] VGG16  [10] 80.89 3.06
I3D  [21] 88.52 2.98
SlowFast  [9] 92.73 3.23
DADA  [23] VGG16  [10] 81.94 2.96
I3D  [21] 89.40 2.92
SlowFast  [9] 95.39 3.08
Table 6: Performance of UString with Global Features extracted from SlowFast network
@@
DoTA [22] UString [5] VGG16 [10] 75.40 2.26
SlowFast [9] 86.38 1.90
DADA [23] UString [5] VGG16 [10] 63.95 1.03
SlowFast [9] 66.14 1.11
We also evaluated the UString framework by integrating global features extracted from the SlowFast network. The results, as presented in Table 6, show a notable improvement in performance compared to the original UString method, which uses VGG-16 global features. Specifically, the Average AP increased to 86.38 on the DoTA dataset and 66.14 on the DADA dataset, while achieving competitive mTTA values of 1.90s and 1.11s, respectively. This demonstrates that incorporating global context through SlowFast features enhances the anticipation capability of the UString model as well.
@@ 4.2.3 Ablation Study
Data LSTM GATv2 SlowFast Avg AP (%) Avg mTTA (s)
DoTA  [22] 83.50 3.14
@@
The ablation results presented in Table 7 on the DoTA and DADA datasets illustrate the individual contribution of each additional component or modification — LSTM, GATv2, and SlowFast — to the overall performance of our accident anticipation framework.
From the results, it is evident that reverting any of the modifications causes a drop in average precision (AP) on both datasets. However, using SlowFast instead of I3D features and/or adding an LSTM network in global feature learning has a much larger impact than the addition of GATv2 in object graph learning. For instance, on the DoTA dataset, the baseline without any of the three modifications achieves an AP of 83.50%. Adding the LSTM alone increases AP considerably to 89.98%, indicating the critical role of temporal modeling. Similarly, replacing I3D with SlowFast as the feature extractor itself yields an AP of 90.03%, highlighting the advantage of powerful spatio-temporal feature representations. Combining LSTM and SlowFast further improves performance to 92.18%, and the best results are achieved when all three components (LSTM, GATv2, and SlowFast) are used together, reaching 92.73% AP and an average mTTA of 3.23s. A similar trend is observed on the DADA dataset. The baseline without any modifications yields 84.04% AP, which is improved to 89.18% by adding the LSTM alone. Replacing I3D with SlowFast features as the only modification results in 92.95% AP. Again, the highest performance with 95.39% AP and 3.08s mTTA is attained when all components are used.
The ablation study also reveals that reverting either modification reduces mTTA. Furthermore, we note that neither of the modifications alone improves the mTTA compared to the baseline Graph(Graph) model.
@@
4.2.4 Analysis of Execution Speed
Table 8 shows the execution time of different stages in the pipeline for different methods. These experiments were conducted on a computing platform with the following configuration: GPU – Tesla P100 with 16 GiB of memory, CPU – 2-core Intel Xeon (at 2.2 GHz), and 29 GiB of RAM.
According to the results, the most time is taken for detecting objects and extracting VGG-16 features from each bounding box (0.063 s/frame), which is common to all comparison methods. In terms of global feature extraction, VGG is the fastest, followed by I3D and SlowFast, respectively. In terms of model inference time (excluding pre-trained feature extractors), STAGNet and Graph(Graph) operate faster than the BNN-based Ustring model. Overall, models using VGG-16 or I3D global features operate close to 10 FPS, while our STAGNet model with SlowFast features is slightly slower at 9 FPS. However, we note that the implementations of neither the STAGNet nor other models have been optimized for efficiency. In fact, none of the comparison methods used in this work have previously reported run-times. Therefore, we expect that these models can be made more efficient. Especially, the time-consuming object feature extraction can be parallelized to improve speed.
Table 8: Execution times and speed of different methods. Times for different stages are reported in seconds per frame (s/f), averaged over DoTA and DADA datasets. The overall speed is reported as frames per second (FPS).
Method Object Detection and Feature Extraction (s/f) Global Feature Extraction (s/f) Inference (s/f) Overall speed (FPS)

### v2

@@ 4.1.2 Networks and Baselines
We use the pretrained YOLO [wang2024yolov10realtimeendtoendobject] and VGG16 [simonyan2014very] models to detect objects and extract object-level features, respectively. Specifically, we utilize the top-19 objects detected, each represented by a feature vector of dimension $d_{1}=4096$. To encode object labels, we apply GloVe word embeddings [pennington2014glove] using the spaCy library, resulting in embedding vectors of size $d_{2}=300$. When setting temporal edges in the object graphs, each object node is connected only to its counterpart in the previous frame ($u=1$). For frame-level graphs, we set the number of temporal neighbors to $k=20$ [thakur2024graph]. The same architecture and configuration, illustrated in Figure 2, are used across all datasets.
For global visual feature extraction, we adopt the SlowFast network [Feichtenhofer2018SlowFastNF], which produces a $h$-dimensional feature vector for each frame, where $h=2304$. All experiments are implemented using the PyTorch framework.
For our re-implementation of AAT-DA [kumamoto2025aat], we used the model parameters presented in the original paper, as follows. The same feature extraction pipeline that is commonly used in previous methods [bao2020uncertainty, karim2022dynamic] was used: Top-19 objects detected were considered, and 4096-dimensional object-level and global features were extracted using VGG16. For all the transformers, the hidden layer dimension and the number of attention heads were set to 1024 and 8, respectively. Dropout rates of 0.3, 0.3, 0.1, and 0.5 were applied to the spatial transformer, object self-attention layer, temporal transformer, and the fully connected layer, respectively.
On the DAD dataset, we compare our approach against several accident anticipation methods: DSA [chan2017anticipating], adaLEA [suzuki2018anticipating], UString [bao2020uncertainty], DSTA [karim2022dynamic], EGSTL [mahmood2023new], AAT-DA [kumamoto2025aat], and Graph(Graph) [thakur2024graph]. These algorithms were selected based on their relevance to this research and widespread use as baselines in similar work. The results for these methods are directly taken from the respective publications. The results for our proposed STAGNet model are based on the same training/testing splits of DAD as those used by these comparison methods.
On the DoTA and DADA datasets, we compared our model against UString [bao2020uncertainty], UString (Geometric), Graph(Graph) [thakur2024graph], and the recent transformer-based model AAT-DA [kumamoto2025aat]. We were unable to compare with the remaining methods because their implementations were not available for us to run on the DoTA and DADA datasets. Note that we implemented UString (Geometric) as an attempt to approximate GSTL [mahmood2023new] by replacing the VGG16-based object-level features in UString [bao2020uncertainty] with geometric features derived from the detected bounding boxes. Furthermore, although Mahmood et al. [mahmood2023new] used a similar number of videos from the DoTA dataset, their selected subset of accident videos differs from ours. We specifically selected ego-involved accidents, whereas they did not. Thus, the DoTA results reported in [mahmood2023new] are not comparable to those in our study.
@@
Our STAGNet model was implemented in PyTorch version 2.0.0 and trained for 20 Epochs using the Adam optimizer with a batch size of 1 and a learning rate of 0.0001. These same hyper-parameters, which are inherited from [bao2020uncertainty, thakur2024graph], were applied to all three datasets.
@@ 4.1.3 Evaluation Metrics
Accident anticipation models are expected to be both accurate and early in their predictions. Consistent with prior work [bao2020uncertainty, mahmood2023new, thakur2024graph, fang2023vision], we evaluate our model using Average Precision (AP) and mean Time-to-Accident (mTTA). AP reflects the correctness of predictions by measuring how well the model distinguishes between frames leading to an accident and those without an accident in the near future. For a given frame, if the predicted probability $p_{t}$ of a future accident exceeds a threshold $\alpha$, it is counted as a positive prediction. By varying $\alpha$, a precision-recall curve is generated, and AP is computed as the area under this curve. A higher AP indicates more accurate predictions.
To determine how early the model can anticipate an accident, we use the Time-to-Accident (TTA) metric, which measures the time difference between the actual onset of an accident and the earliest frame for which $p_{t}>\alpha$. By evaluating TTA across different thresholds $\alpha$, we compute the mean TTA (mTTA) as the average of all corresponding values.
@@ 4.2.1 Comparison with State-of-the-art Methods
Table 2 shows the results of the experiments on DAD. Our proposed algorithm achieves the highest AP and mTTA, closely followed by Graph(Graph) [thakur2024graph]. These two methods perform considerably better than the other comparison methods. The transformer-based method, AAT-ODA [kumamoto2025aat], achieves a very competitive AP but at the expense of a lower mTTA.
Table 2: Performance comparison on the dash-cam accident dataset (DAD)
@@
DAD  [chan2017anticipating] DSA [chan2017anticipating] 48.10 1.34
adaLEA [suzuki2018anticipating] 52.30 3.43
UString [bao2020uncertainty] 53.70 3.53
DSTA [karim2022dynamic] 56.1 3.66
EGSTL [mahmood2023new] 56.30 3.23
Graph(Graph)  [thakur2024graph] 62.80 4.29
AAT-DA  [kumamoto2025aat] 64.0 2.87
STAGNet-Lite (Ours) 50.47 4.24
STAGNet (Ours) 64.36 4.32
As shown in Table 3, for the DoTA and DADA datasets containing ego-involved accidents, our model consistently predicts accidents more precisely and earlier than the comparison methods. On DoTA, the UString (Geometrics) and UString models achieve average precisions of 63.16% and 75.40%, respectively, while the Graph(Graph) model reaches 83.50%. We also evaluated the recent transformer-based method AAT-DA [kumamoto2025aat], which achieved an AP of 72.45% [^3] Note that the implementation of AAT-DA was not made available by the authors. Therefore, we implemented it based on the details given in the paper to the best of our knowledge. For the DAD dataset, we were able to achieve very similar results (63% AP) to those reported in their work (64% AP), suggesting a reasonable re-implementation.. In comparison, our STAGNet model achieves an average precision of 92.73% and an mTTA of 3.23 seconds. Our pruned model, STAGNet-Lite, also achieves comparable results, with an average precision of 91.89% and an mTTA of 3.13 seconds. On the DADA dataset, our model achieves 95.39% AP, outperforming UString (Geometrics) at 56.75%, UString at 63.95%, AAT-DA at 69.36%, and Graph(Graph) at 84.04%. The corresponding mTTA of 3.08 seconds also confirms our methods’ ability to anticipate accidents earlier. STAGNet-Lite achieves an average precision of 94.42% and an mTTA of 3.11 seconds.
@@
DoTA  [yao2022dota] UString (Geometrics)  [mahmood2023new] 63.16 2.14
UString  [bao2020uncertainty] 75.40 2.26
Graph(Graph) [thakur2024graph] 83.50 3.14
AAT-DA* [kumamoto2025aat] 72.45 3.28
STAGNet-Lite (Ours) 91.89 3.13
@@
DADA  [fang2019dada] UString (Geometrics)  [mahmood2023new] 56.75 1.25
UString  [bao2020uncertainty] 63.95 1.03
Graph(Graph)  [thakur2024graph] 84.04 2.92
AAT-DA* [kumamoto2025aat] 69.36 1.36
STAGNet-Lite (Ours) 94.42 3.11
@@
* Re-implemented based on the details given in the original paper.
@@ 4.2.2 Cross-dataset Evaluation
To assess the ability of our model to generalize, we conducted cross-dataset experiments by training on one dataset and evaluating it on another. During this process, any hyperparameter tuning was confined to the training dataset. As shown in Table 4, our method outperforms the comparison methods across both datasets. When trained on the DoTA dataset and tested on DADA, our method achieves an AP of 67.47%, outperforming UString (Geometrics) at 46.39%, UString at 48.26%, AAT-DA at 50.27%, and Graph(Graph) at 61.34%. Additionally, it achieves a higher mTTA of 3.28 seconds, indicating better anticipation of accidents. Similarly, when trained on DADA and evaluated on DoTA, our method achieves an AP of 80.15%, surpassing UString (Geometrics) at 54.85%, UString at 55.63%, AAT-DA at 66.23%, and Graph at 68.55%. It also records a higher mTTA of 3.13 seconds. These results highlight the better generalization ability of our model across datasets.
However, the cross-dataset performance of all tested algorithms is considerably worse than the cross-validation results on the same datasets. To investigate possible reasons for this drop in performance, we probed for differences in video quality and the distribution of object categories involved in the accidents between the DoTA and DADA datasets.
Figure 3 reveals that there is no significant difference between the bit-rates of DoTA and DADA videos. However, there is a noticeable difference in visual quality between the two datasets. Figure 4 shows a few samples. Videos in DADA use the MPEG-4 Visual Simple Profile (SP) codec, whereas those in DoTA use Xvid. Even at similar bit-rates, the Xvid codec may use MPEG-4 Advanced Simple Profile (ASP) features and produce higher quality videos [^4] https://labs.xvid.com/project/. This difference in quality may contribute to the drop in cross-dataset performance. In particular, the higher quality of DoTA videos can produce clearer object boundaries and smoother motion, which can enable the model to learn more detailed spatial-temporal features. However, these learned features may not generalize well to the lower-quality videos in DADA.
As shown in Figure 5(a), the DoTA dataset has very few accidents involving cyclists, road obstacles, pedestrians, and motorbikes compared to DADA. We hypothesized that when trained on DoTA and tested on DADA, these categories would have worse recall than others. However, Figure 5(b) shows that although the recall is low for cyclists and motorbikes, the recall for pedestrians and obstacles is comparatively high despite their low presence in the training data. Therefore, the under-representation of some object categories does not fully explain the drop in cross-dataset performance.
@@
DoTA  [yao2022dota] DADA (bAP = 40.8) UString (Geometrics)  [mahmood2023new] 46.39 1.71
UString  [bao2020uncertainty] 48.26 1.95
Graph(Graph) [thakur2024graph] 61.34 3.12
AAT-DA [kumamoto2025aat] 50.27 3.24
@@
DADA  [fang2019dada] DoTA (bAP = 50.0) UString (Geometrics)  [mahmood2023new] 54.85 2.45
UString  [bao2020uncertainty] 55.63 2.17
Graph(Graph) [thakur2024graph] 68.55 2.86
AAT-DA [kumamoto2025aat] 66.23 2.18
@@
Figure 3: Box plots showing the bit-rates of DoTA and DADA videos.
Figure 4: Sample frames from DADA (top row) and DoTA (bottom row). The visual quality of $DA$ videos is lower than that of DoTA videos.
(a) Distribution of crash-object categories.
(b) Cross-dataset recall for different crash-object categories (at an overall recall of 80%).
Figure 5: (a) Distribution of crash-object categories and (b) category-wise cross-dataset recall for DoTA and DADA datasets. For each test dataset, the recall values are reported at the probability threshold that yields an overall recall of 80%.
4.2.3 Impact of Global Features
We analyze the impact of various global feature extraction methods on our accident anticipation framework. Since spatial features extracted from pretrained image-based networks such as VGG16 focus solely on individual frames, they lack the ability to capture the temporal dynamics across frames. To overcome this limitation, spatio-temporal features extracted using I3D and SlowFast networks were explored. As shown in Table 5, the I3D model improves performance over VGG16. Furthermore, our proposed use of SlowFast yields the best performance, outperforming both VGG16 and I3D. This improvement is attributed to the dual-pathway design of SlowFast, which effectively captures both fast motion cues and slow-evolving spatial context, providing a more comprehensive spatio-temporal representation.
Note that we have kept the object detector (YOLO), object-level feature extractor (VGG16), and the global feature extractor (SlowFast) frozen at their pre-trained weights during training. All prior works known to us have also frozen their object detectors and feature extractors. This is reasonable for the YOLO and VGG16 models, as they have been trained on large-scale datasets containing vehicles, pedestrians, and other road objects. However, the SlowFast networks have been trained on human activity datasets [Feichtenhofer2018SlowFastNF]. Therefore, finetuning them on road traffic data may improve the models performance in accident anticipation. To evaluate this, we finetuned the SlowFast networks on the corresponding training sets of DoTA and DADA. However, there was only a marginal gain in AP and no improvement in mTTA, as evident from the results in Table 5.
Table 5: Comparison of the effects of different global feature extractors on our STAGNet model.
@@ 4.2.3 Impact of Global Features
DoTA  [yao2022dota] VGG16  [simonyan2014very] 80.89 3.06
I3D  [carreira2017quo] 88.52 2.98
SlowFast  [Feichtenhofer2018SlowFastNF] 92.73 3.23
SlowFast (finetuned) 93.24 3.19
DADA  [fang2019dada] VGG16  [simonyan2014very] 81.94 2.96
I3D  [carreira2017quo] 89.40 2.92
SlowFast  [Feichtenhofer2018SlowFastNF] 95.39 3.08
SlowFast (finetuned) 95.81 3.01
Table 6: Performance of UString with global features extracted from SlowFast networks.
@@
DoTA [yao2022dota] UString [bao2020uncertainty] VGG16 [simonyan2014very] 75.40 2.26
SlowFast [Feichtenhofer2018SlowFastNF] 86.38 1.90
DADA [fang2019dada] UString [bao2020uncertainty] VGG16 [simonyan2014very] 63.95 1.03
SlowFast [Feichtenhofer2018SlowFastNF] 66.14 1.11
We also evaluated the UString framework by integrating global features extracted from the SlowFast networks. The results, as presented in Table 6, show a notable improvement in performance compared to the original UString method, which used VGG16 global features. This demonstrates that incorporating global context through SlowFast features enhances the UString model as well.
@@ 4.2.4 Ablation Study
Data LSTM GATv2 SlowFast AP (%) mTTA (s)
DoTA  [yao2022dota] 83.50 3.14
@@
The ablation results presented in Table 7 demonstrate the individual contributions of each additional component or modification — LSTM, GATv2, and SlowFast — to the overall performance of our accident anticipation framework.
From the results, it is evident that reverting any of the modifications causes a drop in average precision (AP) on both datasets. However, using SlowFast instead of I3D features and/or adding an LSTM network in global feature learning has a much larger impact than the addition of GATv2 in object graph learning. For instance, on the DoTA dataset, the baseline without any of the three modifications achieves an AP of 83.50%. Adding the LSTM alone increases AP considerably to 89.98%, indicating the critical role of temporal modeling. Similarly, replacing I3D with SlowFast as the feature extractor itself yields an AP of 90.03%, highlighting the advantage of powerful spatio-temporal feature representations. Combining LSTM and SlowFast further improves performance to 92.18%, and the best results are achieved when all three components (LSTM, GATv2, and SlowFast) are used together, reaching 92.73% AP and an average mTTA of 3.23 seconds. A similar trend is observed with the DADA dataset. The baseline without any modifications yields 84.04% AP, which is improved to 89.18% by adding the LSTM alone. Replacing I3D with SlowFast features as the only modification results in 92.95% AP. Again, the highest performance, with 95.39% AP and 3.08 s mTTA, is attained when all components are used.
The ablation study also reveals that reverting either modification reduces mTTA. Furthermore, we note that neither modification, when applied in isolation, improves the mTTA compared to the baseline Graph(Graph) model.
@@
4.2.5 Analysis of Computational Efficiency and Execution Speed
Tables 8 and 9 show the execution times and the number of floating-point operations (FLOPs), respectively, for the different stages of the pipeline across different methods. These experiments were conducted on a computing platform with the following configuration: GPU – Tesla P100 with 16 GiB of memory, CPU – 2-core Intel Xeon (at 2.2 GHz), and 29 GiB of RAM.
According to the results, object detection and VGG16 feature extraction from bounding boxes (19 most confident detections) account for the majority of the computational cost: 0.063 s/frame, 322.46 GFLOPs/frame. This is a burden common to all compared methods, except for our pruned variant, STAGNet-Lite. In terms of global feature extraction, VGG is the fastest, followed by I3D and SlowFast, respectively. In terms of model inference time (excluding pre-trained feature extractors), STAGNet and Graph(Graph) operate faster than the BNN-based Ustring model and the transformer-based AAT-DA model. Overall, UString and Graph(Graph) models using either VGG16 or I3D global features operate at close to 10 FPS, while our STAGNet model with SlowFast features is slightly slower at 9 FPS. However, we note that the implementations of neither STAGNet nor other models have been optimized for efficiency. In fact, none of the comparison methods used in this work have previously reported run-times. Therefore, we expect that these models could be further optimized. Especially, the time-consuming object feature extraction can be parallelized to improve speed.
On the other hand, our pruned model, STAGNet-Lite, has achieved 21 FPS, which is much faster than all other models. It also maintains high AP and mTTA values, falling slightly short of the full STAGNet model but higher than the remaining models for ego-involved accidents (DoTA and DADA datasets). However, it shows a more prominent drop in AP compared to the full model for non-ego-involved accidents (DAD dataset), which is expected due to pruning the modules responsible for object detection and object-object interaction modeling. Nevertheless, STAGNet-Lite can be practically useful for resource-constrained ADAS environments due to its execution speed and good performance in anticipating ego-involved accidents, which is the most important class of accidents.
Table 8: Execution times and speed of different methods. Times for different stages are reported in seconds per frame (s/f), averaged over DoTA and DADA datasets. The overall inference speed is reported as frames per second (FPS). Note that the inference time of AAT-DA shows the time taken to generate its driver attention maps as well.
Method Object Detection & Feature Extraction (s/f) Global Feature Extraction (s/f) Inference (s/f) Overall speed (FPS)